The centralized employee Travel Department of Johnson Company has expenses of $364,800. The department has serviced a total of 19,200 travel reservations for the period. The Northeast Division has made 14,400 reservations during the period, and the Pacific Division has made 4,800 reservations.

How much should each division be charged for travel services? Do not round interim calculations.

Northeast Division $
Pacific Division $

Campbell Company has income from operations of $53,768, invested assets of $188,000, and sales of $488,800. Use the DuPont formula to compute the rate of return on investment. If required, round your answers to two decimal place.

a. Profit margin %
b. Investment turnover
c. Rate of return on investment %

The Consumer Division of Peanut Co. has income from operations of $72,600 and assets of $275,000. The minimum acceptable rate of return on assets is 8%.

What is the residual income for the division?
$
Answer #1

a) Expense allocated

Northeast division = 364800/19200*14400 = $273600

Pacific division = 364800/19200*4800 = 91200

b) Profit margin = Operating income/Sales = 53768/488800 = 11%

Investment turnover = 488800/188000 = 2.6 Times

Return on investment = 11*2.6 = 28.6%

c) Residual income = 72600-(275000*8%) = $50600